OpenStudy (aayushi.somani):

Integrate : (x-1) root(1-x-x^2)

myininaya (myininaya):

Have you tried completing the square inside the square root first?

myininaya (myininaya):

This might end up being a trig sub.

myininaya (myininaya):

i will be back if you have questions

OpenStudy (aayushi.somani):

yeah i tried it.. but that didnt worked !

OpenStudy (anonymous):

u can take (x-1) as u and the rest term as v and integrate by parts

OpenStudy (aayushi.somani):

@ikram002p please help !

OpenStudy (ikram002p):

addy :o is it like this ? \(\int (x-1) \sqrt (1-x-x^2)\)

OpenStudy (aayushi.somani):

yeah exactly!

myininaya (myininaya):

The trig sub seems to work Not saying there aren't other ways though

myininaya (myininaya):

@aayushi.somani may I ask how far you got with that method?

myininaya (myininaya):

to what integral have you transformed it?

OpenStudy (aayushi.somani):

i could not figure out anything

myininaya (myininaya):

can you tell me what substitutions you performed ?

myininaya (myininaya):

some people like to do two here i just like one

myininaya (myininaya):

When you say you could not figure out anything does that mean you did not do anything with the trig sub even though you said you did?

myininaya (myininaya):

I'm confused

OpenStudy (aayushi.somani):

i tried getting a trig sub but unfortunatly i couldnt.. thats what it is.. @ikram002p could u figure out?

myininaya (myininaya):

did you complete the square inside the square root part?

OpenStudy (aayushi.somani):

if i could have done it then y would i ask this ques..??? @myininaya i asked so that anyone could help me..

myininaya (myininaya):

Oh I'm sorry I didn't know you couldn't do the algebra either

myininaya (myininaya):

So by nothing means absolutely nothing I get it

OpenStudy (aayushi.somani):

from what i did m getting a wrong answer so consider i did nothing and now please help me from the very begnning..

myininaya (myininaya):

\[ax^2+bx+c=a(x^2+\frac{b}{a}x)+c=a(x^2+\frac{b}{a}x+(\frac{b}{a})^2)+c-a(\frac{b}{a})^2 \\ =a(x+\frac{b}{a})^2+c-a \frac{b^2}{a^2}=a(x+\frac{b}{a})^2+c-\frac{b^2}{a}\] I hope this helps you complete the square Please follow these steps I prefer you ask me questions on how to do this than just using the end result there

OpenStudy (ikram002p):

 i dont know if this work next to use tring mmm sin maybe ?